Sol Summary – August 11th

 In Sol Summary

Jon Clarke

Sol 23 (or 103)

Today dawned slightly foggy but it soon cleared up to be a nice sunny day, though still with a stiff breeze. Good EVA weather! In the morning we fetched water to replenish our tanks for the remaining three people who are having their weekly shower tonight.

Today’s EVA went only a short distance into the crater, but lasted over four hours and achieved many goals. Yusuke and Anushree were suited, Anastasiya road shotgun. First Yusuke carried out a 3D scan of a field of sorted stone polygons, the same site trenched by Paul four sols ago. Aunshree collected hypolith data while he was going this. When this was done there was a quick diversion to an area of impact melt rocks for sampling. The team then made their way back to Lake Cornell where they collected further hypolith data. Their final tasks, back at the hab, was to sample the pink biofilms at the micro oasis just below the ridge crest.

Tonight I made Palak Paneer and rice. We had an extended discussion afterwards to discuss how we will close out the station when we depart next week.

Tomorrow will be a very busy sol. We are aiming to do two EVAs because of the expected bad weather coming on Sunday may preclude EVAs for the following couple of days. Until then it’s goodnight from me and goodnight from the rest of the crew.

Start typing and press Enter to search